undeadknight26 (undeadknight26):

3. Simplify the given expression to radical form and justify each step by identifying the properties of rational exponents used. All work must be shown. (x^(2/3))/(x^(4/9))

sammixboo (sammixboo):

Frick

sammixboo (sammixboo):

I forgot this

undeadknight26 (undeadknight26):

I dunno what to do from: (x^(2/3))/(x^(4/9)) = x^((2/3)-(4/9)) = x^(2/6)

sammixboo (sammixboo):

Tell me how you got there. Maybe it will boost my memory ahaha

undeadknight26 (undeadknight26):

When you divide you subtract the exponents right? Thats what i did.

undeadknight26 (undeadknight26):

@kawaiicat123

sammixboo (sammixboo):

Omg I totally forgot

sammixboo (sammixboo):

My mind is blank

undeadknight26 (undeadknight26):

Ish okay miss sammy you helped me out enough already XD

sammixboo (sammixboo):

I will still try ahahaha

undeadknight26 (undeadknight26):

@YanaSidlinskiy

OpenStudy (yanasidlinskiy):

Gimme some time;)

sammixboo (sammixboo):

Thanks for helping him because I couldnt >_< x_x

OpenStudy (yanasidlinskiy):

Is this how it's supposed to look like??? \[\frac{ x^(\frac{ 2 }{ 3 }) }{x^\frac{ 4 }{ ?9} }\]

OpenStudy (yanasidlinskiy):

should be 4/9 on the bottom

undeadknight26 (undeadknight26):

no parentheses..

sammixboo (sammixboo):

Yeah, no parentheses, that is just to show it is separate from everything else

OpenStudy (yanasidlinskiy):

Kk!!! First you have to subtract the exponents..\[\frac{ 2 }{ 3 } - \frac{ 4 }{ 9 }\] Wut do u get?xD!!

sammixboo (sammixboo):

I thought that is what you do omg

sammixboo (sammixboo):

xD

sammixboo (sammixboo):

It just came back to me, so I will help check and all

OpenStudy (yanasidlinskiy):

:) Thats wut u gotta doxD!!

undeadknight26 (undeadknight26):

x^2/6?

OpenStudy (yanasidlinskiy):

*Close* Not quite. Give it another shot? Maybe?

undeadknight26 (undeadknight26):

Im out of ammo o.o

OpenStudy (yanasidlinskiy):

Hhaha!!! Kk!! It's x^2/9 *because.....you have to rewrite the first problem by multiplying top and bottom to get 6/9. Then you have to subtract 2/9.

sammixboo (sammixboo):

Yeppers

OpenStudy (yanasidlinskiy):

\[\frac{ 2 }{ 3 } \times \frac{ 3 }{ 3 } = \frac{ 6 }{ 9 } - \frac{ 4 }{ 9 } = \frac{ 2 }{ 9 }\] to get* 2/9

OpenStudy (yanasidlinskiy):

Undastood, Knight?!!

OpenStudy (yanasidlinskiy):

Kk..I guess u dun wanna answerxD!! Anyways....Next is next..Gimme a secxD!!

sammixboo (sammixboo):

Ahaha I think he is doing something but I will grab him real quick brb

undeadknight26 (undeadknight26):

Sorry mates! Yesh i understand banana!

OpenStudy (yanasidlinskiy):

Hhahaha!!! No no!! It's fine;)!! To put it in radical form......this is how it's gonna look like:) \[x^\frac{ 2 }{ 9 } = \sqrt[9]{x^2}\]

OpenStudy (yanasidlinskiy):

And..I guess that's all:)

undeadknight26 (undeadknight26):

OHHH alright thanks!

OpenStudy (yanasidlinskiy):

Ya!!!!!:) Welcome!!:) Blew ur mind again. Right;)!! Lol!! I guess it's good to have a friend that helps with Algebra!!;)xD!!

sammixboo (sammixboo):

xD Yeah